OpenStudy (lena772):

find two positive numbers that satisfy the given requirements. The sum of the first number squared and the second is 27 and the product is a maximum

OpenStudy (lena772):

\[x^2+y=27\]

OpenStudy (lena772):

@agent0smith @Hero @phi

hero (hero):

So y = 27 - x^2 set x = 0, to find y

OpenStudy (lena772):

y=27

OpenStudy (anonymous):

\[product P=x^2y=x^2(27-x^2)=27x^2-x^4\] \[find \frac{ dP }{dx } and~ put =0,find ~the~values~of~x\] then find\[\frac{ d^2P }{dx^2 } at this point\]